Question: Line 1 has y intercept (0,5) and x intercept (4,0) Find the equation of line 1 in each case and then write it in standard form with integral coefficients. This is what I have, m=5-0=-5/4 slope then x1+x2/2=2-4/2=3.
y1+y2/2=0+4=2 x=3 y=2 . Now I don't know what to do. Please help

Answer by stanbon(75887)

Question: Line 1 has y intercept (0,5) and x intercept (4,0) Find the equation of line 1 in each case and then write it in standard form with integral coefficients. This is what I have, m=5-0=-5/4 slope then x1+x2/2=2-4/2=3.
y1+y2/2=0+4=2 x=3 y=2
-------------
m=-5/4 and the line passes thru (4,0)
Use the point-slope formula to get:
y-0=(-5/4)(x-4)
y=(-5/4)x+5
To convert to standard form multiply thru by 4 to get:
4y=-5x+20
5x+4y=20
------------
Cheers,
Stan H.